Find the equation of the line passing through (-1,-2) and parallel to 6x+2y=5
Answer by likaaka(51)

The slopes of parallel lines are the same so first you must find the slope of the line 6x+2y=5. I would change that equation to slope-intercept form or y=mx+b where m is your slope and b is your x-intercept
6x+2y=5, subtract 6x from both sides
6x+2y - 6x=5 -6x
2y=-6x+5, divide both side by 2
(2y)/2=(-6x+5)2
y=((-6x)/2)+(5/2)
y=-3x+(5/2)
According to slope-intercept form, you now know that the slope is -3
Given the slope and a point (-1,-2), you can use point-slope form to get the equation of the line
Point-slope form is (y-ysub1)=m(x-xsub1) where ysub1 and xsub1 is the given point and m is the slope
(y-(-2))=-3(x-(-1))
y+2=-3(x+1), I'm not sure in what form you must give your solution but if it is accepted in point slope form then that's it but I would get rid of the parenthesis on the right side of the equation by distributing the -3
y+2=-3x-3, If you must answer in slope-intercept form then solve for y
y+2 -2=-3x-3 -2
y=-3x-5